# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v3

@@ 3.4.3 Tracking
Similarly to detection tasks, DMs for tracking tasks are used as downstream processes by concatenating the diffusion head to the feature extraction backbone. However, a spatial-temporal fusion module has been added to the diffusion head to exploit temporal video features[194]. DiffusionTrack [195] localizes the target in a progressive diffusion manner, which is claimed to better handle challenging scenarios. The method in [196] exploits spatial-temporal weighting to suppress the probability of the tracker changing the target to the distractors. It, however, reports under-performance compared to MixFormer.
@@
Bridging the gap between digital and physical realms, 3D reconstruction and rendering are integral to various creative technologies. In film and animation, they enable the creation of detailed digital models that blend seamlessly with live-action footage. Video games and digital twins leverage these technologies for dynamic environmental rendering. VR and AR use 3D reconstruction to create immersive and interactive experiences, with AR integrating digital content into real-world contexts. With recent AI technologies, 3D reconstruction and rendering have become faster and closer to reality. In particular, neural radiance fields and Gaussian Splatting enable artists and film producers to create shots that cannot be one in the real shooting environments.
@@ 3.5.1 Depth Estimation
As deep learning can capture monocular cues such as object size, texture gradients, and perspective, depth estimation from a single image can produce accurate results. There have been attempts to use transformers, such as [199] and [200], and diffusion models, such as [205] and [207]. Amongst these, Depth Anything v2 [202] has become a state-of-the-art monocular depth estimation method. It is built on the previous version [201], jointly trained on large-scale labeled and unlabeled images and uses semantic priors from pretrained encoders. Depth Anything v2 significantly outperforms V1 in fine-grained details and robustness by using synthetic images and pseudo-labeled real images, as well as by extracting intermediate features from DINOv2 [166], which is trained with vision transformers. One of the notable capabilities of Depth Anything v2 is its ability to predict depth of transparent and reflective surfaces.
@@ 3.5.2 Neural Radiance Fields
Since their introduction, there have been many variants of NeRFs aimed at improving their performance. Mip-NeRF360 [204] proposed unbounded anti-aliased technique achieving full 360 degree content. Google Research [210] trains NeRF from noisy RAW images captured in the dark scene, allowing changing viewpoint, focus, exposure, and tone mapping simultaneously. With segmentation techniques significantly advanced (see Section 3.4.1), there have been integrations utilizing semantic segmentation to enhance 3D representation [212]. DSEM-NeR [203] integrates the pretrained CLIP model to extract multimodal features—including color, depth, and semantics—from multi-view 2D images, thereby enhancing the reconstruction quality of complex scenes.
@@
The initial application of NeRFs to dynamic scenes was undertaken by Pumarola et al. [208], known as D-NeRF. However, the current leading method for generating high-quality novel views of real dynamic scenes is TiNeuVox [211]. It enhances temporal information by interpolating voxel features before feeding them into the radiance network to estimate density and color, similar to ordinary NeRF. DynVideo-E [214] adds an MLP to predict motion fields but focuses on human-centric content. PaReNeRF [217] addresses large-scale dynamic scenes using patch-based sampling. The main drawback of these methods is the large model size and/or long training time. Therefore, $K$-planes [218] propose a simple planar factorization for volumetric rendering, achieving low memory usage (1000$\times$ compression over a full 4D grid). Wavelet transform are employed in [215] to further reduce model size. KFD-NeRF [216] incorporates a Kalman filter-guided deformation field for more accurate motion estimation.
@@ 3.6.1 Image Compression
Since the first neural image codec [329] was proposed in 2016, numerous learning-based image compression methods have been developed, with significant performance improvements reported [330, 331]. Driven by the latest advances in neural network architectures, neural image codecs now outperform standard image codecs. Instead of using CNNs as the basic network structure, transformer-based architectures have become popular, offering the potential for better compression efficiency. Notable examples include SwinT-ChARM [226], STF [227] and LIC-TCM [228]. SwinT-ChARM [226] employs Swin transformers for non-linear transforms and outperforms the latest standard image codec, the Versatile Video Coding (VVC) Test Model (VTM, All Intra). STF [227] is based on a symmetrical transformer framework containing absolute transformer blocks in both the down-sampling encoder and the up-sampling decoder, which also shows improved rate-quality performance over VTM. LIC-TCM [228] exploits the local modeling ability of CNN and the non-local modeling performance of transformers, and proposes a parallel transformer-CNN mixture block. This new network structure, together with a channel-wise entropy model based on attention modules using Swin transformers, contributes to the superior performance of STF, with a more than 10% bitrate saving over VTM.
An alternative approach to learned image coding is based on advanced generative models. Early works [332, 333] employed GANs to generate more photo realistic results with improved visual quality. Although these models fail to outperform conventional, CNN-based or transformer-based approaches, when distortion-based quality metrics, e.g., PSNR, are used for performance evaluation, they have been reported to perform well when perceptual quality models, such as MS-SSIM [334] and VMAF [335], or subjective tests are employed to measure perceived video quality. More recently, diffusion models have been applied in image compression to allow realistic reconstruction at ultra-low bitrates [229] achieving competitive performance compared to GAN-based models [230]. However, it should be noted that some of these generative models aim to generate (or synthesize) images with “perfect realism” rather than reconstruct results which are most similar to the original content. Notable work in this category includes image codecs using score-based generative models [231] and the diffusion-based residual augmentation codec (DIRAC) [232]. Moreover, another type of generative model based on INR has been employed for image compression; this learns a mapping between the spatial coordinates and the respective pixel values for the input image. The learned INR model is then compressed through parameter quantization and model compression to minimize the required bitrate. Notable INR-based image codecs include COIN/COIN++ [234, 235] and [236] that combine SIREN networks [233] with positional encoding.
@@ 3.6.2 Video Compression
i) The enhancement of conventional coding tools focuses on employing deep learning techniques to improve the performance of one (or multiple) coding modules in a standard-applicant codec. These modules include intra prediction [341], inter prediction [342], in-loop filtering [343], post filtering [344] and resolution re-sampling [345]. To facilitate efficient integration, the MPEG Joint Video Experts Team (JVET) built a test model in 2022 based on VTM 11, named Neural Network-based Video Coding (NNVC) [346], with its latest version NNVC-7.1 containing two major learning-based coding tools, neural-network based intra prediction and in-loop filtering, which has achieved an up to 13% coding gain over VTM 11 (Random Access mode) [347]. However, this learning-based codec requires much higher computational complexity (up to 477 kMACs/pixel) and high-spec GPU support compared to conventional codecs. Meanwhile, members of the Alliance of Open Media (AOM) have also developed multiple CNN-based coding tools for the next generation of video coding standard beyond AV1. The latest proposals focus on the trade-off between performance and complexity, with one of them based on inloop filtering and super-resolution, which achieves an average BD-rate saving of 3.9% (in PSNR) over AVM, the test model of AV2, but only requires a much lower computational complexity (below 1.5kMACs/pixel) [348]. More recently, research has been conducted to further improve the performance of these learning-based coding tools utilizing more advanced network architectures, including ViTs [349], and diffusion models [239]. There are also investigations on applying preprocessing before compression [350, 351], where the training of the deep preprocessors is based on proxy video codecs and/or rate-distortion loss functions to simulate the behavior of conventional video coding algorithms.
@@
ii) End-to-end optimized neural video codecs. Alongside the enhancement of coding tools in conventional video codecs, more recent research activities have focused on using neural networks to implement the whole coding workflow, enabling data-driven end-to-end optimization. The performance of these neural video codecs has advanced significantly in the last five years, since the first attempt, DVC [352], was published. DVC matched the performance of a fast implementation of H.264 (x264). However currently, learned video coding algorithms (e.g., DCVC-FM [353] and DCVC-LCG [354]) are able to compete or even outperform the state-of-the-art standard codecs, such as VVC VTM under certain coding configurations. These learning-based methods often focus on enhancement from different perspectives, including feature space conditional coding (e.g., FVC [355] and DCVC [356]), instance adaptation  [357, 358], and motion estimation (e.g., DCVC-DC [359]). New architectures have also been proposed such as CANF-VC [360] based on a video generative model, MTMT [237] using a masked image modeling transformer-based entropy model and VCT [238] based on a video compression transformer. It is noted that although promising coding performance has been achieved in the aforementioned works, these neural video codecs (in particular those based on autoencoder backbones) are typically associated with high computational complexity (especially in the decoder), which constrains their deployment for practical applications. To address this issue, researchers attempted to achieve complexity reduction while maintaining the coding performance through model pruning and knowledge distillation [361, 362].
@@ 3.6.3 Audio Compression
Similarly to images and videos, learning-based solutions have also been researched to compress audio signals, and most neural audio codecs are based on VQ-VAE [265]. SoundStream [368] is one of such models, which can encode audio content at various bitrates. It is based on a residual vector quantizer (RVQ) which trades off between rate, distortion, and complexity. This work has been further enhanced with a multi-scale spectrogram adversary and a loss balancer mechanism, resulting in improved rate-distortion performance. A more advanced universal model has been further developed [369] based on improved adversarial and reconstruction losses, which can compress different types of audio. RVQ has also been extended from a single scale to multiple scales [370], which performs hierarchical quantization at variable frame rates.
@@ 3.7 Visual Quality Assessment
In recent years, quality assessment methods have been enhanced using deep learning techniques. The resulting learning-based quality models can quickly adapt to a specific type of content, leading to better performance compared to conventional, hand-crafted quality metrics. This section provides a brief summary of existing works in this research area, and highlights the main challenges which should be addressed in the near future. A more comprehensive overview of the image and video quality assessment literature can be found in [373, 374, 375].
@@ 3.7.1 Quality assessment models
Image and video quality assessment methods can be classified into two primary categories according to the availability of the corresponding reference content to the distorted test version: full-reference and no-reference models [^55] Reduced-reference quality metrics do existing in the literature, but the research in this field is less active in recent years.. Prior to the AI era, conventional visual quality methods often exploit different characteristics of the human vision system and capture relevant information related to structural similarity (such as in SSIM and its variants [376, 377, 378]), distortion [379, 380, 381], and artifacts [382, 383, 384]. In many cases, the extracted features are further processed by models that simulate texture masking [385], contrast sensitivity [386], and saliency [387]. These hand-crafted quality models have also been combined with features within a regression-based framework in order to achieve more accurate prediction performance - VMAF is one such example [335]. When neural networks are involved for feature extraction, they are trained to capture information which can directly contribute to quality prediction through an end-to-end optimization strategy. Initially, convolutional neural networks were typically used, with notable examples such as DeepQA [388], LPIPS [389] and CONTRIQUE [390] for image quality assessment, and TLVQA [391], C3DVQA [392] and DeepVQA [393] for video quality assessment. Recent works have been reported to achieve better performance when Vision Transformers (ViTs) (or similar variants) are employed due to the effectiveness of their self-attention mechanism. Important works in this class include IQT [248], TRes [249], SaTQA [250], FastVQA [251] and RankDVQA [252]. The former has been further extended as DOVER [253] and COVER [254] when aesthetic and/or semantic aspects in the content are taken into account.
@@
To support the training and validation of learning-based quality assessment models, image or video databases containing ground-truth subjective quality scores are typically employed. Commonly used image quality databases include LIVE [404], CSIQ [380], TID2013 [405], PieAPP and PIPAL, while video quality databases such as LIVE-VQA [406], KoNViD-1K [407], YouTube UGC [408] and LIVE-VQC [409] are typically employed for benchmarking in the literature. There are also databases developed that investigate the impact of specific video formats and/or artifacts, such as LIVE-YT-HFR [410] focusing on frame rates, VSR-QAD [411] on spatial resolution (or super-resolution artifacts), BAND-2k [412] on banding artifacts and Maxwell [403]/BVI-Artifact [413] containing multiple artifacts commonly produced in video streaming. Based on these databases, many learning-based quality assessment models are trained to minimize the difference (L1 or L2 norm) between predicted quality indices and subjective scores. However, due to the limited number of ground-truth quality labels associated with these databases and the resourcing-costing nature for collecting subjective data through human participants involved in psychophysical experiments, this type of training methodology cannot offer satisfactory performance, in particular when the model capacity is large. Moreover, since the experimental settings and conditions used for quality labeling are different in these databases, intra-database cross-validation is always required due to the limited model generalization and potential overfitting problems.
To address these issues, various proxy quality metrics have been used to label images and videos, which avoid expensive subjective tests and enable the generation of a large amount of training material with pseudo-ground-truth quality annotations. To further improve the reliability of quality labels, instead of learning the absolute values of the quality labels, ranking-inspired training strategies have been developed, which focus on improving the monotonicity characteristics of quality. Important examples based on these weakly supervised training methodologies include RankIQA [414] and UNIQUE [415] for the image quality assessment task, and VFIPS [416] and RankDVQA [252] for video quality assessment. Moreover, different self-supervised learning approaches have also been employed, which transform quality labeling to an auxiliary task. For example, CONTRIQUE [390] learns relevant features from an unannotated image database based on the prediction of distortion types and degrees through contrastive learning. This method has been further applied to video quality assessment, resulting in a contrastive video quality estimator, CONVIQT [417]. More recently, quality-aware contrastive loss has been designed in [418, 255] to stabilize the learning process.

### v4

@@ 3.4.3 Tracking
Similarly to detection tasks, DMs for tracking tasks are used as downstream processes by concatenating the diffusion head to the feature extraction backbone. However, a spatial-temporal fusion module has been added to the diffusion head to exploit temporal video features[194]. DiffusionTrack [195] localizes the target in a progressive diffusion manner, which is claimed to better handle challenging scenarios. The method in [196] exploits spatial-temporal weighting to suppress the probability of the tracker changing the target to the distractors. It, however, reports underperformance compared to MixFormer.
@@
Bridging the gap between digital and physical realms, 3D reconstruction and rendering are integral to various creative technologies. In film and animation, they enable the creation of detailed digital models that blend seamlessly with live-action footage. Video games and digital twins leverage these technologies for dynamic environmental rendering. VR and AR use 3D reconstruction to create immersive and interactive experiences, with AR integrating digital content into real-world contexts. With recent AI technologies, 3D reconstruction and rendering have become faster and closer to reality. In particular, neural radiance fields and Gaussian Splatting enable artists and film producers to create shots that cannot be done in the real shooting environments.
@@ 3.5.1 Depth Estimation
As deep learning can capture monocular cues such as object size, texture gradients, and perspective, depth estimation from a single image can produce accurate results. There have been attempts to use transformers, such as [199] and [200], and diffusion models, such as [205] and [207]. Amongst these, Depth Anything v2 [202] has become a state-of-the-art monocular depth estimation method. It is built on the previous version [201], jointly trained on large-scale labeled and unlabeled images and uses semantic priors from pretrained encoders. Depth Anything v2 significantly outperforms V1 in fine-grained details and robustness by using synthetic images and pseudo-labeled real images, as well as by extracting intermediate features from DINOv2 [166], which is trained with vision transformers. One of the notable capabilities of Depth Anything v2 is its ability to predict the depth of transparent and reflective surfaces.
@@ 3.5.2 Neural Radiance Fields
Since their introduction, there have been many variants of NeRFs aimed at improving their performance. Mip-NeRF360 [204] proposed an unbounded anti-aliased technique achieving full 360 degree content. Google Research [210] trains NeRF from noisy RAW images captured in the dark scene, allowing changing viewpoint, focus, exposure, and tone mapping simultaneously. With segmentation techniques significantly advanced (see Section 3.4.1), there have been integrations utilizing semantic segmentation to enhance 3D representation [212]. DSEM-NeR [203] integrates the pretrained CLIP model to extract multimodal features—including color, depth, and semantics—from multi-view 2D images, thereby enhancing the reconstruction quality of complex scenes.
@@
The initial application of NeRFs to dynamic scenes was undertaken by Pumarola et al. [208], known as D-NeRF. However, the current leading method for generating high-quality novel views of real dynamic scenes is TiNeuVox [211]. It enhances temporal information by interpolating voxel features before feeding them into the radiance network to estimate density and color, similar to ordinary NeRF. DynVideo-E [214] adds an MLP to predict motion fields but focuses on human-centric content. PaReNeRF [217] addresses large-scale dynamic scenes using patch-based sampling. The main drawback of these methods is the large model size and/or long training time. Therefore, $K$-planes [218] propose a simple planar factorization for volumetric rendering, achieving low memory usage (1000$\times$ compression over a full 4D grid). Wavelet transforms are employed in [215] to further reduce model size. KFD-NeRF [216] incorporates a Kalman filter-guided deformation field for more accurate motion estimation.
@@ 3.6.1 Image Compression
Since the first neural image codec [329] was proposed in 2016, numerous learning-based image compression methods have been developed, with significant performance improvements reported [330, 331]. Driven by the latest advances in neural network architectures, neural image codecs now outperform standard image codecs. Instead of using CNNs as the basic network structure, transformer-based architectures have become popular, offering the potential for better compression efficiency. Notable examples include SwinT-ChARM [226], STF [227] and LIC-TCM [228]. SwinT-ChARM [226] employs Swin transformers for non-linear transforms and outperforms the latest standard image codec, the Versatile Video Coding (VVC) Test Model (VTM, All Intra). STF [227] is based on a symmetrical transformer framework containing absolute transformer blocks in both the down-sampling encoder and the up-sampling decoder, which also shows improved rate-quality performance over VTM. LIC-TCM [228] exploits the local modeling ability of CNN and the non-local modeling performance of transformers, and proposes a parallel transformer-CNN mixture block. This new network structure, together with a channel-wise entropy model based on attention modules using Swin transformers, contributes to the superior performance of STF, with more than 10% bitrate savings over VTM.
An alternative approach to learned image coding is based on advanced generative models. Early works [332, 333] employed GANs to generate more photo-realistic results with improved visual quality. Although these models fail to outperform conventional, CNN-based or transformer-based approaches, when distortion-based quality metrics, e.g., PSNR, are used for performance evaluation, they have been reported to perform well when perceptual quality models, such as MS-SSIM [334] and VMAF [335], or subjective tests are employed to measure perceived video quality. More recently, diffusion models have been applied in image compression to allow realistic reconstruction at ultra-low bitrates [229] achieving competitive performance compared to GAN-based models [230]. However, it should be noted that some of these generative models aim to generate (or synthesize) images with “perfect realism” rather than reconstruct results that are most similar to the original content. Notable work in this category includes image codecs using score-based generative models [231] and the diffusion-based residual augmentation codec (DIRAC) [232]. Moreover, another type of generative model based on INR has been employed for image compression; this learns a mapping between the spatial coordinates and the respective pixel values for the input image. The learned INR model is then compressed through parameter quantization and model compression to minimize the required bitrate. Notable INR-based image codecs include COIN/COIN++ [234, 235] and [236] that combine SIREN networks [233] with positional encoding.
@@ 3.6.2 Video Compression
i) The enhancement of conventional coding tools focuses on employing deep learning techniques to improve the performance of one or more coding modules in a standard-compliant codec. These modules include intra prediction [341], inter prediction [342], in-loop filtering [343], post-decode filtering [344] and resolution re-sampling [345]. To facilitate efficient integration, the MPEG Joint Video Experts Team (JVET) built a test model in 2022 based on VTM 11, named Neural Network-based Video Coding (NNVC) [346], with its latest version NNVC-7.1 containing two major learning-based coding tools, neural-network based intra prediction and in-loop filtering, which has achieved an up to 13% coding gain over VTM 11 (Random Access mode) [347]. However, this learning-based codec requires much higher computational complexity (up to 477 kMACs/pixel) and high-spec GPU support compared to conventional codecs. Meanwhile, members of the Alliance of Open Media (AOM) have also developed multiple CNN-based coding tools for the next generation of video coding standard beyond AV1. The latest proposals focus on the trade-off between performance and complexity, with one of them based on in-loop filtering and super-resolution, which achieves an average BD-rate saving of 3.9% (in PSNR) over AVM, the test model of AV2, but only requires a much lower computational complexity (below 1.5kMACs/pixel) [348]. More recently, research has been conducted to further improve the performance of these learning-based coding tools utilizing more advanced network architectures, including ViTs [349], and diffusion models [239]. There are also investigations on applying preprocessing before compression [350, 351], where the training of the deep preprocessors is based on proxy video codecs and/or rate-distortion loss functions to simulate the behavior of conventional video coding algorithms.
@@
ii) End-to-end optimized neural video codecs. Alongside the enhancement of coding tools in conventional video codecs, more recent research activities have focused on using neural networks to implement the whole coding workflow, enabling data-driven end-to-end optimization. The performance of these neural video codecs has advanced significantly in the last five years, since the first attempt, DVC [352], was published. DVC matched the performance of a fast implementation of H.264 (x264). However currently, learned video coding algorithms (e.g., DCVC-FM [353] and DCVC-LCG [354]) are able to compete or even outperform the state-of-the-art standard codecs, such as VVC VTM under certain coding configurations. These learning-based methods often focus on enhancement from different perspectives, including feature space conditional coding (e.g., FVC [355] and DCVC [356]), instance adaptation  [357, 358], and motion estimation (e.g., DCVC-DC [359]). New architectures have also been proposed such as CANF-VC [360] based on a video generative model, MTMT [237] using a masked image modeling transformer-based entropy model and VCT [238] based on a video compression transformer. It is noted that, although promising coding performance has been achieved by the aforementioned contributions, these neural video codecs (in particular those based on autoencoder backbones) are typically associated with high computational complexity (especially in the decoder), which constrains their deployment for practical applications. To address this issue, researchers are now focused on complexity reduction while maintaining coding performance through model pruning and knowledge distillation [361, 362].
@@ 3.6.3 Audio Compression
Similarly to images and videos, learning-based solutions have also been researched to compress audio signals, and most neural audio codecs are based on VQ-VAE [265]. SoundStream [368] is one such model, which can encode audio content at various bitrates. It is based on a residual vector quantizer (RVQ) which trades off rate, distortion and complexity. This work has been further enhanced with a multi-scale spectrogram adversary and a loss balancer mechanism, resulting in improved rate-distortion performance. A more advanced universal model has been further developed [369] based on improved adversarial and reconstruction losses, which can compress different types of audio. RVQ has also been extended from a single scale to multiple scales [370], which performs hierarchical quantization at variable frame rates.
@@ 3.7 Visual Quality Assessment
In recent years, quality assessment methods have been enhanced using deep learning techniques. The resulting learning-based quality models can quickly adapt to a specific type of content, leading to better performance compared to conventional, hand-crafted quality metrics. This section provides a brief summary of existing work in this research area, and highlights the main challenges which should be addressed in the near future. A more comprehensive overview of the image and video quality assessment literature can be found in [373, 374, 375].
@@ 3.7.1 Quality assessment models
Image and video quality assessment methods can be classified into two primary categories according to the availability of the corresponding reference (un processed) content. These are referred to as full-reference and no-reference models [^55] Reduced-reference quality metrics do exist in the literature, but research in this field has been less active in recent years.. Prior to the AI era, conventional visual quality methods often exploited characteristics of the human vision system capturing information related to structural similarity (such as in SSIM and its variants [376, 377, 378]), distortion [379, 380, 381], and artifacts [382, 383, 384]. In many cases, the extracted features are further processed by models that simulate texture masking [385], contrast sensitivity [386], and saliency [387]. These hand-crafted quality models have also been combined with features within a regression-based framework in order to achieve more accurate prediction performance - VMAF is one such example [335]. When neural networks are used for feature extraction, they are trained to capture information which can directly contribute to quality prediction through an end-to-end optimization strategy. Initially, convolutional neural networks were used for this, with notable examples such as DeepQA [388], LPIPS [389] and CONTRIQUE [390] for image quality assessment, and TLVQA [391], C3DVQA [392] and DeepVQA [393] for video quality assessment. Recent works have been reported to achieve better performance when Vision Transformers (ViTs) (or similar variants) are employed due to the effectiveness of their self-attention mechanism. Important works in this class include IQT [248], TRes [249], SaTQA [250], FastVQA [251] and RankDVQA [252]. The former has been further extended as DOVER [253] and COVER [254] when aesthetic and/or semantic aspects in the content are taken into account.
@@
To support the training and validation of learning-based quality assessment models, image or video databases containing ground-truth subjective quality scores are typically employed. Commonly used image quality databases include LIVE [404], CSIQ [380], TID2013 [405], PieAPP and PIPAL, while video quality databases such as LIVE-VQA [406], KoNViD-1K [407], YouTube UGC [408] and LIVE-VQC [409] are typically employed for benchmarking in the literature. There are also databases developed that investigate the impact of specific video formats and/or artifacts, such as LIVE-YT-HFR [410] focusing on frame rates, VSR-QAD [411] on spatial resolution (or super-resolution artifacts), BAND-2k [412] on banding artifacts and Maxwell [403]/BVI-Artifact [413] containing multiple artifacts commonly produced in video streaming. Based on these databases, many learning-based quality assessment models are trained to minimize the difference (L1 or L2 norm) between predicted quality indices and subjective scores. However, due to the limited number of ground-truth quality labels associated with these databases and the resource requirements associated with collecting subjective data using human participants in psychophysical experiments, this type of training methodology cannot offer satisfactory performance, in particular when the model capacity is large. Moreover, since the experimental settings and conditions used for quality labeling are different in these databases, intra-database cross-validation is always required due to the limited model generalization and potential overfitting problems.
To address these issues, various proxy quality metrics have been used to label images and videos, which avoid expensive subjective tests and enable the generation of a large amounts of training material with pseudo-ground-truth quality annotations. To further improve the reliability of quality labels, instead of learning the absolute values of the quality labels, ranking-inspired training strategies have been developed, which focus on improving the monotonicity characteristics of quality. Important examples based on these weakly supervised training methodologies include RankIQA [414] and UNIQUE [415] for the image quality assessment task, and VFIPS [416] and RankDVQA [252] for video quality assessment. Moreover, different self-supervised learning approaches have also been employed, which transform quality labeling to an auxiliary task. For example, CONTRIQUE [390] learns relevant features from an unannotated image database based on the prediction of distortion types and degrees through contrastive learning. This method has been further applied to video quality assessment, resulting in a contrastive video quality estimator, CONVIQT [417]. More recently, quality-aware contrastive loss has been designed in [418, 255] to stabilize the learning process.